Calculate the quartile deviation and it's co efficient of quartile deviation for the following items.
X: 5 12 9 15 24 18 12
1
Expert's answer
2019-10-07T09:23:48-0400

At first we neet to sort the initial data


"5,9,12,12,15,18,24"

To calculate quartile deviation

"{Q_{dev}} = \frac{{{Q_{3/4}} - {Q_{1/4}}}}{2}"

We need to calculate the 1/4 quartile and 3/4 quartile. Because the number of elements in our list is odd, the 1/4 quartile is defined as the average of the median of the "\frac{{n - 1}}{2}" smallest elements and the median of the "\frac{{n + 1}}{2}" smalles elements. The median of the first 3 smalles elements "5,9,12" is "9" and the median of the first 4 smalles elements "5,9,12,12" is "\frac{{9 + 12}}{2} = 10.5" . Thus "{Q_{1/4}} = \frac{{9 + 10.5}}{2} = 9.75" . The same procedure we shall do for the 3/4 quartile (but we shall use largest instead smallest). I.e. The median of the first 3 largest elements "15,18,24" is "18" and The median of the first 4 largest elements "12,15,18,24" is "\frac{{15 + 18}}{2} = 16.5" . Thus "{Q_{3/4}} = \frac{{16.5+ 18}}{2} = 17.25" . Now we can calculate the quartile deviation


"{Q_{dev}} = \frac{{{Q_{3/4}} - {Q_{1/4}}}}{2} = \frac{{17.25 - 9.75}}{2} = 3.75"

And we can calculate quartive variation coefficient


"{Q_{var }} = \frac{{{Q_{3/4}} - {Q_{1/4}}}}{{{Q_{3/4}} + {Q_{1/4}}}} \cdot 100[\% ] = \frac{5}{{18}} \cdot 100[\% ] \approx 27.778[\% ]"